# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

### v1

@@ IV Constrained Trajectory Potential Games
Therefore, using Theorem 1, finding local GNE trajectories is equivalent to finding local solutions to (5). The conventional approach to solving the optimization in (5) relies on techniques such as inter-point methods [23] or sequential quadratic programming [24]. Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [23], hoping to explore the solution space of local Nash equilibria. However, random exploration using gradient-based methods becomes computationally prohibitive for large $N$ and inefficient as the same local equilibria can be arrived at from multiple random initializations. To address this, we utilize a Bayesian inference perspective in the following section to recover all local minima of (5) in an efficient manner.
V Identifying Local Nash Equilibria via Bayesian Inference
In this section, we propose our approach to solve (5) from a Bayesian inference perspective. We reformulate the optimal control problem (5) into an equivalent Bayesian inference problem by utilizing the reference trajectory and constraint violations as virtual measurements that provide evidence for the optimal inference of the system trajectory in (5). At its core, the Bayesian inference problem is about approximating the probability density function (PDF) of the trajectories given measurement information derived from the objective and constraints in (5). Then, the high probability regions of the PDF to be approximated will correspond to the local GNEs of interest. This enables using the sample efficient implicit particle filter [30], which is designed to sample from the high probability regions of the PDF. The resulting identified equilibria then serve as a warm-start initialization for a downstream optimizer such as IPOPT, which finds the fine-tuned trajectory. We first formulate the MulitNash-PF problem and then outline the algorithm that realizes the framework.
V-A MultiNash-PF framework via Bayesian Inference
Given problem (5), we view the objective function and constraints as a source of optimality information to facilitate the inference of the optimal trajectory. This motivates the introduction of a state-space model that sets us up to construct the Bayesian reformulation as a joint state and control inference via the objective function and constraint as measurements. We construct the virtual system as
$x_{t+1}$ $=f(x_{t},u_{t}),$ $y_{x,t}$ $=x_{t}+v_{t},$
$u_{t+1}$ $=w_{t},$ $y_{g,t}$ $=\psi\left(g\left(x_{t},u_{t}\right)\right)+\eta_{t},$ (7)
for $0\leq t\leq\tau$, where $x_{t}$ and $u_{t}$ are the states of the virtual system, $y_{x,t}$, $y_{g,t}$ are measurements, and $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances. The function $\psi(\cdot)$ represents a barrier function that can capture the degree of constraint violation by a trajectory. We chose the softplus barrier function for its continuity in the forthcoming filtering algorithm. The solfplus function is expressed as:
$\psi(g\left(x_{t},u_{t}\right))=\ln\left(1+\exp(g(x_{t},u_{t}))\right)/\alpha,$ (8)
where $\alpha$ is a tuneable parameter that controls the strictness of enforcing constraints. Considering the virtual system (V-A), we relax the problem (5) using soft constraints as:
$\begin{array}[]{ll}\underset{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}{\min}&\frac{1}{2}\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}\\ &+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},\\ \quad\;\mathrm{s.t.}&\begin{bmatrix}x_{0}\\ u_{0}\end{bmatrix}=\begin{bmatrix}\hat{x}_{0}\\ w_{-1}\end{bmatrix},\begin{bmatrix}x_{t+1}\\ u_{t+1}\end{bmatrix}=\begin{bmatrix}f(x_{t},u_{t})\\ 0_{m}\end{bmatrix}+\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},\vspace{0.2cm}\\ &\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix}=\begin{bmatrix}x_{t}\\ \phi\left(g(x_{t},u_{t})\right)\end{bmatrix}+\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},0\leq t\leq\tau,\end{array}$ (9)
@@ V-A MultiNash-PF framework via Bayesian Inference
where $Q_{\eta}$ is a positive definite weight matrix that assigns the importance of constraint satisfaction. Problem (9) differs from (5) in that it considers a soft inequality constraint using the barrier function (8). Further, (9) is also known as the moving horizon estimation problem [31] which aims to find the minima of the disturbances $\left\{v_{t}^{\star},w_{t}^{\star},\eta_{t}^{\star}\right\}_{t=0}^{\tau}$. A direct attempt at solving (9) using traditional gradient-based methods poses a similar computational burden in identifying local Nash equilibria compared to problem (5). However, the above connection between (5) and the moving horizon estimation problem (9) highlights the potential to treat the problem from a Bayesian inference perspective. It can be seen that the constraints in (9) are the same as the considered virtual system in (V-A), suggesting that state inference can be performed via efficient estimation methods. To proceed, we rewrite (V-A) compactly as
@@
$\bar{v}_{t}=\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},\;\bar{f}($ $\bar{x}_{t})=\begin{bmatrix}f(x_{t},u_{t})\\ {0}_{m}\end{bmatrix},\;\bar{h}(\bar{x}_{t})=\begin{bmatrix}x_{t}\\ \psi(g\left(x_{t},u_{t}\right))\end{bmatrix}.$
Here, we relax $\bar{w}_{t}$ and $\bar{v}_{t}$ for $0\leq t\leq\tau$ to be stochastic disturbances modeled as normally distributed random variables with PDF
$\bar{v}_{t}\sim\mathcal{N}(0_{n+c},\bar{Q}),\;\bar{v}_{\tau}\sim\mathcal{N}(0_{n+c},\bar{Q}_{\tau}),\;\bar{w}_{t}\sim\mathcal{N}(0_{n+m},\bar{R}),$ (11)
with $\bar{Q}\coloneqq\mathop{\rm blkdiag}(Q^{-1},Q_{\eta}^{-1})$, $\bar{Q}_{\tau}\coloneqq\mathop{\rm blkdiag}(Q_{\tau}^{-1},Q_{\eta}^{-1})$, $\bar{R}\coloneqq\mathop{\rm blkdiag}(0_{n\times n},R^{-1})$. The state estimation relevant to (10) is to estimate the state $\bar{x}_{t}$ given the measurements $\bar{y}_{t}$. We must set the virtual measurement’s value to correspond to the objective of the optimization in (5). We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$. While $y_{g,t}$ is to satisfy constraints through (8). Further, we observe that the barrier function in (8) outputs zero when the constraints are satisfied. Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
@@
The Bayesian inference problem pertaining to (10) is to characterize the conditional PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. In the following, lemma (1) elucidates that characterizing the maximum a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ holds equivalence to solving (5).
@@
We refer the interested reader to [12, Thm. 1] for the proof. The above lemma implies that the local GNEs in (5) translate to multiple local maximums of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Hence, the Bayesian inference problem for (10) involves characterizing the multi-modal PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Generally, state estimation of a multi-modal PDF does not admit closed-form solutions [32]. Hence, we reside to approximate solutions. A powerful approach is to use Monte Carlo methods that empirically approximate the distribution via a set of $J$ particles as
@@
where $\delta(\cdot)$ is the dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$. After identifying the local GNE through the above approximation, we extract the trajectories that appear as local GNE and use them to initialize an optimization solver. This completes the exposition of the MultiNash-PF framework.
We employ the implicit particle filter  [30] due to its two-fold benefit for implementing the MultiNash-PF framework. First, as suggested in lemma 1, the local maximums (i.e high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs. Hence, focusing our sampling efforts towards these high-probability regions can reduce the required samples for the approximation in (12). Second, we want to avoid the traditionally needed resampling step in particle filtering methods to ensure that the inferred trajectories are smooth. Fortunately, the implicit particle filter satisfies both of the aforementioned criteria. We leverage the Markovian property of the virtual system (10) to approximate (12). Using the Bayes’ rule, we have
@@
for $0\leq t\leq\tau$. This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can recursively obtain $\bar{x}_{t}^{j}$, reducing the complexity involved in directly sampling from a high-dimensional PDF. As shown in [30], we use implicit importance sampling and identify an implicit mapping to sample from high probability regions of (13). For this purpose, we make a local Gaussian assumption around $\bar{x}_{t-1}^{j}$ such that
$p(\bar{x}_{t}^{j}|\bar{y}_{0:t})\sim\mathcal{N}\left(\hat{m}_{t}^{j},\hat{\Sigma}_{t}^{x,j}\right),$ (14)
where $\hat{m}_{t}^{j}$ and $\hat{\Sigma}_{t}^{x,j}$ are the mean and error covariance of $\bar{x}_{t}^{j}$, respectively. These quantities are then found recursively by using a nonlinear Kalman filter. Then, the implicit map $X(\gamma_{t}^{j})$ can be computed in closed form as
$\bar{x}_{t}^{j}=\hat{m}_{t}^{j}+\sqrt{\hat{\Sigma}_{t}^{x,j}}\gamma_{t}^{j},$ (15)
where $\gamma_{t}^{j}$ is a sample from $\mathcal{N}(0,I_{n})$ and $I_{n}$ denotes the identity matrix of size $n$.
@@
5:for $t=0,1,\ldots,\tau-1$ do
6:     for $i=1,2,\ldots,J$ do
7:Run Kalman prediction step:
8:         $\left(m_{t+1}^{j},\Sigma^{x,j}_{t+1}\right)\leftarrow\text{{\bf UT}}(\bar{x}_{t}^{j},\Sigma_{t}^{x,j},\bar{R},\bar{f})$
9:Run Kalman update step:
10:         if $t\leq\tau-2$ then
11:              $(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1},\Sigma^{xy,j}_{t+1})\leftarrow\text{{\bf UT}}(m_{t+1}^{j},\Sigma_{t+1}^{x,j},\bar{Q},\bar{h})$
12:         else
13:              $(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1},\Sigma^{xy,j}_{t+1})\leftarrow\text{{\bf UT}}(m_{t+1}^{j},\Sigma_{t+1}^{x,j},\bar{Q}_{\tau},\bar{h})$
14:         end if
15:         $K_{t+1}^{j}\leftarrow\Sigma_{t+1}^{xy,j}(\Sigma_{t+1}^{y,j})^{-1}$
16:         $\hat{m}^{x,j}_{t+1}=m_{t+1}^{j}+K_{t+1}^{j}\left(\bar{y}_{t+1}-\hat{\bar{y}}_{t+1}^{j}\right)$
17:         $\hat{\Sigma}_{t+1}^{x,j}\leftarrow\Sigma_{t+1}^{x,j}-K_{t+1}^{j}\Sigma^{y,j}_{t+1}(K_{t+1}^{j})^{\top}$
18:Draw a high-probability sample $\gamma_{t+1}^{j}\sim\mathcal{N}(0,I_{n})$ for the implicit mapping in (15):
19:         $\bar{x}_{t+1}^{j}\leftarrow\hat{m}_{t+1}^{x,j}+\sqrt{\hat{\Sigma}_{t+1}^{x,j}}\gamma_{t+1}^{j}$
20:Weight and normalize samples:
21:         $w_{t+1}^{j}=\frac{w_{t}^{j}p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})}{\sum_{j=1}^{J}w_{t}^{j}p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})},\qquad p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})\sim\mathcal{N}(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1})$
22:     end for
23:     Perform resampling if necessary
24:end for
25:Extract trajectories from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
V-B Implicit Particle Filter via Unscented Kalman filter
As in [30], we note that the local Gaussian approximation in (14) can be recursively computed through a nonlinear Kalman filter evolving according to (13) for all $j=1,\ldots,J$. Our nonlinear Kalman filter of choice is the unscented Kalman filter due to its high filtering accuracy for low-dimensional state space models [33]. The unscented Kalman filter is based on the unscented transform, which utilizes deterministic samples, known as sigma points, to trace the evolution of the mean and covariance of a Gaussian random variable through a nonlinear function. The unscented Kalman filter [33] utilizes the unscented transform in the Kalman prediction and measurement update steps to keep track of the Gaussian statistics through the state transition and measurement function in (10). From now on, we use UT to denote Unscented Transform algorithm.
Using the unscented Kalman filter, we can recursively compute the local Gaussian approximation (14) at time $t$. Then, we implicitly map the particle to the high-probability regions by sampling $\gamma_{t}^{i}\sim\mathcal{N}(0,I_{n})$ for (15) to obtain the particle $\bar{x}_{t}^{i}$. If necessary, we perform resampling for the particle trajectory $\bar{x}_{0:t}^{i}$. This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs. This completes the unscented implicit particle filter, which has identified the coarse estimates of multiple local Nash equilibria in (5). As the last step, the estimates of local equilibria are extracted and used as warm start initialization to the IPOPT optimizer to find the optimal local GNEs. To extract the equilibria, we carefully observe the trajectories and select the trajectories that are repeatedly taken by multiple particles. The pragmatic extraction of equilibria from the particle trajectories is nontrivial and is left for future study. The implementation details of the proposed MultiNash-PF through the unscented implicit particle filter are provided in Algorithm 1.
VI Numerical Simulations

### v2

@@ III-A Local Generalized Nash Equilibrium Trajectories
The conventional approach to solving the optimization (1) relies on techniques such as interior point methods [24] or sequential quadratic programming [25]. Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [24], hoping to explore the solution space of local Nash equilibria. However, randomly exploring the nonconvex optimization landscape via gradient-based methods is computationally prohibitive and inefficient, as the same local equilibria can be arrived at from multiple random initializations. To address this, we utilize a Bayesian inference approach in the following section to efficiently recover all local minima of (1).
IV Identifying Local Nash Equilibria via Bayesian Inference
In this section, we propose our approach to solve (1) from a Bayesian inference perspective. We reformulate the optimal control problem (1) into an equivalent Bayesian inference problem by utilizing the reference trajectory and constraint violations as virtual measurements that provide evidence for the optimal inference of the system trajectory in (1). To this end, we first reformulate problem (1) by considering soft-constraints and introducing auxiliary variables $y_{x,t}$ and $y_{g,t}$ as:
$\min_{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}$ $\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},$
$\mathrm{s.t}\quad$ $\begin{bmatrix}x_{t+1}\\ u_{t+1}\end{bmatrix}=\begin{bmatrix}f(x_{t},u_{t})\\ 0_{m}\end{bmatrix}+\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},$
$\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix}=\begin{bmatrix}x_{t}\\ \psi\left(g(x_{t},u_{t})\right)\end{bmatrix}+\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},$
$\begin{bmatrix}x_{0}\\ u_{0}\end{bmatrix}=\begin{bmatrix}\hat{x}_{0}\\ w_{t-1}\end{bmatrix},\quad 0\leq t\leq\tau,$ (7)
where $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances, and $Q_{\eta}\in\mathbb{R}^{n_{c}\times n_{c}}$ assigns the importance of constraint satisfaction. Further, $\psi$ is a barrier function used to quantify the degree of constraint violation by defining
$\psi\left(g(x_{t},u_{t})\right)=\frac{1}{\alpha}\ln\left(1+\exp\left(g\left(x_{t},u_{t}\right)\right)\right),$ (8)
where $\alpha$ controls the strictness enforcing constraints.
@@ IV Identifying Local Nash Equilibria via Bayesian Inference
We note that the problem in (IV) resembles a moving horizon estimation problem [32]. A direct attempt at solving (IV) using traditional gradient-based methods poses a similar computational burden in identifying all of the local equilibria compared to problem (1). However, the above connection between (1) and the moving horizon estimation problem (IV) highlights the potential to treat the problem from a Bayesian inference perspective. To this end, we construct the following virtual state-space model from (IV) as
@@
$\bar{v}_{t}=\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},\;\bar{f}($ $\bar{x}_{t})=\begin{bmatrix}f(x_{t},u_{t})\\ {0}_{m}\end{bmatrix},\;\bar{h}(\bar{x}_{t})=\begin{bmatrix}x_{t}\\ \psi(g\left(x_{t},u_{t}\right))\end{bmatrix},$
where we relax $\bar{w}_{t}$ and $\bar{v}_{t}$ for $0\leq t\leq\tau$ to be stochastic disturbances modeled as Gaussian-distributed random variables with density
$\bar{v}_{t}\sim\mathcal{N}(0_{n+c},\bar{Q}),\;\bar{v}_{\tau}\sim\mathcal{N}(0_{n+c},\bar{Q}_{\tau}),\;\bar{w}_{t}\sim\mathcal{N}(0_{n+m},\bar{R}),$
with $\bar{Q}\coloneqq\mathop{\rm blkdiag}(Q^{-1},Q_{\eta}^{-1})$, $\bar{Q}_{\tau}\coloneqq\mathop{\rm blkdiag}(Q_{\tau}^{-1},Q_{\eta}^{-1})$, $\bar{R}\coloneqq\mathop{\rm blkdiag}(0_{n\times n},R^{-1})$. The state estimation relevant to (9) is to estimate the state $\bar{x}_{t}$ given the measurements $\bar{y}_{t}$. The value of the virtual measurements must be set in a way that steers the inference problem to have the same optima as in problem (1). We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$. While $y_{g,t}$ is to satisfy constraints through (8). Observing that the barrier function in (8) outputs zero when the constraints are satisfied, we set the virtual measurement for the barrier function equal to $0_{c}$. Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
@@
The Bayesian inference problem pertaining to (9) is one that characterize the conditional distribution $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. In the following, lemma (1) elucidates that equivalence holds between a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ and the problem in (1).
@@
We refer the interested reader to [16, Thm. 1] for the proof. The above lemma implies that the local GNEs in (1) translate to multiple local maximums of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Hence, the Bayesian inference problem for (9) involves characterizing the multi-modal PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Generally, state estimation of a multi-modal PDF does not admit closed-form solutions [33]. Hence, we resort to approximate solutions. A powerful approach is to use Monte Carlo methods that empirically approximate the distribution via a set of $J$ particles as
@@
where $\delta(\cdot)$ is the dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$.
To achieve the approximation in (10), we employ the implicit particle filter  [31] due to the twofold benefit it holds to implement the MultiNash-PF framework. First, as suggested in lemma 1, the local maximums (i.e., high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs. Hence, focusing our sampling efforts towards these high-probability regions can reduce the number of required samples for the approximation in (10). This approach aligns precisely with the principle of implicit importance sampling, whose objective is to build a mapping hat draws particles from the high-probability regions of the filtering distribution in (10). Second, we want to avoid the traditionally needed resampling step in particle filtering methods to ensure that the inferred trajectories are smooth. Fortunately, the implicit particle filter does not suffer from the problem of particle degeneracy, and filtering can be achieved by disregarding resampling.
We leverage the Markovian property of the virtual system (9) to recursively approximate (10). Using Bayes’ rule, we have
@@
for $0\leq t\leq\tau$. This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can obtain $\bar{x}_{t}^{j}$. As shown in [31], we use implicit importamapping sampling and identify an implicit mapping to sample from high probability regions of (11). The implicit map is constructed by considering sampling from a reference random vector $\xi_{t}$ with density $p(\gamma_{t})$ and map it to high probability regions of $p(x_{0:\tau}\,|\,y_{0:\tau})$, such that a sample $\gamma^{i}_{t}$ is mapped to $x_{t}^{i}$. The mapping connects the highly probable regions of $p(\xi)$ to $p(x_{0:\tau}\,|\,y_{0:\tau})$ by letting
$F(x_{t}^{i})-\min F(x_{t}^{i})=G(\gamma_{t}^{i})-\min G(\gamma_{t}^{i}),$ (12)
where $F(x_{t}^{i})=-\log\left(p(x_{t}^{i}\,|\,y_{1:t})\right)$ and $G(\gamma_{t}^{i})=-\log\left(p(\gamma_{t}^{i})\right)$. Finding an explicit solution to (12) is computationally intractable due to the involved nonconvex optimization. To bypass this computational challenge, a local Gaussian approximation around $\bar{x}_{t-1}^{j}$ is employed such that
$p(\bar{x}_{t}^{j}|\bar{y}_{0:t})\sim\mathcal{N}\left(\hat{m}_{t}^{j},\hat{\Sigma}_{t}^{x,j}\right),$
where $\hat{m}_{t}^{j}$ and $\hat{\Sigma}_{t}^{x,j}$ are the mean and covariance of $\bar{x}_{t}^{j}$, respectively. These quantities can be approximated recursively using an Unscented Kalman filter (UKF) [34] for all $j=1,\ldots,N$. Then, the implicit map can be computed in closed form as
$\bar{x}_{t}^{j}=\hat{m}_{t}^{j}+\sqrt{\hat{\Sigma}_{t}^{x,j}}\gamma_{t}^{j},$
where $\gamma_{t}^{j}$ is a sample from $p(\gamma_{t})\sim\mathcal{N}(0,I_{n})$ and $I_{n}$ denotes the identity matrix of size $n$.
This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs. This completes the unscented implicit particle filter, identifying the coarse estimates of multiple local Nash equilibria in (1).
In the last step, the local equilibria are extracted to be from $\{x_{0:\tau}\}_{j=1}^{J}$. As viewed in Fig. 4, the particle set contains multiple trajectories for each GNE. Therefore, to isolate the GNE trajectories, a clustering method is used to aggregate the particles close to each other in one cluster. One can leverage methods such as hierarchical clustering [35] or DBSCAN [36]. It should be noted that clustering methods require a metric to compute distances between the data points. The Fréchet distance [37] is utilized as our metric to compute distances between the trajectories. The clustering of trajectories will result in one cluster for each GNE by taking the average of trajectories over each cluster and feed them as a warm start to the IPOPT solver to obtain the local GNEs. The algorithmic steps of the proposed MultiNash-PF are provided in Algorithm 1.
@@
5:Employ Algorithm-1 from [31] combined with UKF to obtain coarse solutions, $\{x_{0:\tau}\}_{j=1}^{J}$.
6:Extract trajectories through clustering from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
V Numerical Simulations